Consider the following probability distribution.

x 2 4 9
p(x) 1/4 1/4 1/2

a.) Find the sampling distribution for the same mean x (sample statistic) for a random sample of n=3 measurements from this distribution. Put the answers in ascending order for x (sample statistic).

x
p(x)

Is x an unbiased estimator of the mean?

b.) Find the sampling distribution of the sample median m for a random sample of n=3 measurements from this distribution. Put the answers in ascending order for m.

M
p(M)
Answer #1

number of samples =3*3*3 =27

thefefore probability of each sample =1/27

a)

x 2.00 2.67 3.33 4.00 4.33 5.00 5.67 6.67 7.33 9.00
P(x)    1/27    1/9    1/9    1/27    1/9    2/9    1/9    1/9    1/9    1/27

b)

x 2.00 4.00 9.00
P(x)    7/27 13/27    7/27